# Demonstration of Efficient Radon Removal by Silver-Zeolite in a Dark Matter Detector

### v2

@@ II.3 Expected decay rates
Campaign 1 0.252 $\pm$ 0.002 0.254 $\pm$ 0.001111Rate calculated from constant fit after the reference time.
0.251 $\pm$ 0.009222Rate calculated at the reference time.
Campaign 2 0.262 $\pm$ 0.002 0.260 $\pm$ 0.00111footnotemark: 1
0.262 $\pm$ 0.01122footnotemark: 2
Campaign 3 0.266 $\pm$ 0.001
14.5 $\pm$ 0.0822footnotemark: 2
@@
1.07 $\pm$ 0.0222footnotemark: 2
Table S1: Rate in phase I (background) and phase IV (trap open) with either the quality or 214Po cut for all campaigns. The rate in phase IV at the reference time is also included for all campaigns.
Figure S3: Top panel Expected decay rates with either the quality cut (light green curves) or 214Po cut (dark green curves) applied, derived from modeling of the phase III data from each campaign (green histograms), as well as the phase IV data (red histograms). Bottom panel The 90% confidence level (CL) intervals of the inverse R-values over time, obtained with the quality cut (red curves) and 214Po cut (50%) applied (blue curves). The vertical dashed line defines the reference time where the extrapolated rate with quality cut (light green line) is equal to 76.5 Hz.
The rates in phase I with the quality cut are nearly consistent, with campaigns 2 and 3 each exhibiting a slightly higher rate than the previous campaign. This small increase may be attributed to the buildup of 210Po on the inner surface or volume of the SPC due to consecutive injections of radon over time. Based on the expected buildup and the lifetime of 210Po, a small increase in rate—on the order of 0.01 Hz—is expected, which aligns well with the observed data. Similarly, for the 214Po cut, the rates align closely within the uncertainties with a slight increase, likely due to the presence of additional 210Po, contributing to more mis-reconstructed events. The same trend is observed in phase IV when comparing the rates of campaigns 1 and 2 with the quality and 214Po cuts. Nonetheless, the ratio between the rates in phase IV and phase I for campaigns 1 and 2 approaches 1 (see Fig. 3 in the main text), indicating that the radon rate quickly reduces to pre-injection levels in campaigns 1 and 2. This also demonstrates that the performance of the radon trap aligns well with the achievable sensitivity of our experimental setup.
IV S4. R-values
In this section, we describe how the R-value (specifically $R(t)^{-1}$, see Eq. 1 in the main text) is obtained for each measurement campaign. As mentioned in the main text, the observed count rate in phase IV is fit with a generic model consisting of a exponential function plus a constant component, in order to smooth–over statistical fluctuations and remove dependence on the binning of the counts over time. This fit was done with an MCMC following the same procedure followed for the expected rate trend described in section S2. Separate results with and without the 214Po cut are derived by first defining a log–likelihood function for the phase IV model of each campaign, at central time of each bin $i$, scaling the phase III decay rate trend by $R(t_{i})^{-1}$ and allowing the background rate $\eta$ to vary as a nuisance parameter:
$\log\mathcal{L}_{R}(i)\ =\$$h_{4}(i)\log\nu(t_{i})-\nu(t_{i})-\frac{1}{2}\ \left(\frac{\eta-\eta_{\mathrm{obs}}}{\sigma_{\eta}}\right)^{2},$ (S3)
where the latter term is a Gaussian constraint term for the background rate with respect to the observed background rate (from phase I) and its uncertainty ($\eta_{\mathrm{obs}}$ and $\sigma_{\eta}$, respectively). $h_{4}(i)$ is the observed count rate model for phase IV in time bin $i$. The expected count rate in the center of each time bin, $\nu(t_{i})$, is equal to the expected rate trend from phase III (see Eq. S1) multiplied by $R(t_{i})^{-1}$, plus $\eta$. For the calculation of $R(t_{i})^{-1}$ with 214Po cuts, we conservatively assume a 50% detection efficiency for 214Po events, due to the possibility that many isotopes may have settled on the detector surface and then emit alphas away from the detector volume.
A $90\%$ confidence level (CL) interval is calculated for each time bin using the Feldman-Cousins method [Feldman:1997qc]. This is repeated for 500 randomly–drawn MCMC samples from the fit of the phase III rate model and the phase IV observed rate model to include statistical uncertainty for both fits. This yields an ensemble of slightly–varying 90% CL intervals for $R(t_{i})^{-1}$ for each time bin, of which we take the median result for each.
The $R(t)^{-1}$ 90% CL intervals are shown in the bottom panels of Fig. S3 for each campaign. From these results, we determine the final R-value, $R(t)$, at the reference time, which is denoted by vertical dashed lines in Fig. S3. The derived values and their associated 90% CL intervals for both cuts are summarized in Table S2.
Quality cut 214Po cut (50%)
Campaign 1 [4006.9, inf] [5684.9, 28337]
[…]
Campaign 3 [5.4086, 5.5079] [11.395, 11.886]
Table S2: 90% CL intervals (lower and upper limits) of R-values at the reference time with either quality cut or 214Po cut (50%) applied for each campaign. The 90% lower confidence limit (LCL) is reported as the final result for each campaign and shown in Fig. 4 in the main text.

### v3

@@ 4.1 Rate comparison
Campaign 1 0.252 $\pm$ 0.002 0.254 $\pm$ 0.001 0.251 $\pm$ 0.009
Campaign 2 0.262 $\pm$ 0.002 0.260 $\pm$ 0.001 0.262 $\pm$ 0.011
Campaign 3 0.266 $\pm$ 0.001 14.5 $\pm$ 0.08
@@
  • a
    Rate calculated from constant fit after the reference time.
  • b
    Rate calculated at the reference time.
Figure 5: Ratio between the rate in phase IV (trap open) at the reference time and the rate in phase I (background run) for each campaign with the trap filled with silver-zeolite (campaigns 1 and 2) and activated charcoal (campaign 3). The blue circle and red square represent the rate ratio obtained with either just the quality cut or with 214Po cut applied, respectively. Complete radon removal correspond to a rate ratio of 1.
From Table 1, the rates in phase I with the quality cut are nearly consistent, with campaigns 2 and 3 each exhibiting a slightly higher rate than the previous campaign. This small increase may be attributed to the buildup of 210Po on the inner surface or volume of the SPC due to consecutive injections of radon over time. Based on the expected buildup and the lifetime of 210Po, a small increase in rate—on the order of 0.01 Hz—is expected, which aligns well with the observed data. Similarly, for the 214Po cut, the rates align closely within the uncertainties with a slight increase, likely due to the presence of additional 210Po, contributing to more mis-reconstructed events. The same trend is observed in phase IV when comparing the rates of campaigns 1 and 2 with the quality and 214Po cuts. Nonetheless, the ratio between the rates in phase IV and phase I for campaigns 1 and 2 approaches 1 (see Fig. 5), indicating that the radon rate quickly reduces to pre-injection levels in campaigns 1 and 2. This also demonstrates that the performance of the radon trap aligns well with the achievable sensitivity of our experimental setup.
4.2 Radon reduction ratio calculation
To further quantify the performance of the trap, we calculate the radon reduction ratio $R$ (or R-value) from phase IV, defined as
$R(t)=\frac{R_{\text{exp}}(t)}{R_{\text{obs}}(t)}.$ (3)
Here, $R_{\text{exp}}(t)$ is the expected rate obtained from the extrapolated rate trend (see Sec. 3.3) and $R_{\text{obs}}(t)$ is the observed rate from the experimental data after the trap is opened in phase IV. The latter is fit with a generic model consisting of an exponential function plus a constant component, in order to smooth–over statistical fluctuations and remove dependence on the binning of the counts over time. This fit was done with an MCMC following the same procedure to obtain the expected rate trend described in Sec. $3.$
The R-value at a given time in phase IV is determined by optimizing a Poisson likelihood (with a Nelder-Mead algorithm Virtanen:2019joe ; Gao:2012guu ) comparing the observed count model with the expected count rate scaled by $R(t)^{-1}$. Specifically, separate results with and without the 214Po cut are obtained with a log-likelihood function for the phase IV model of each campaign, at the central time of each bin $i$, scaling the phase III decay rate trend by $R(t_{i})^{-1}$ and allowing the background rate $\eta$ to vary as a nuisance parameter, such as
$\log\mathcal{L}_{R}(i)\ =\$$h_{4}(i)\log\nu(t_{i})-\nu(t_{i})-\frac{1}{2}\ \left(\frac{\eta-\eta_{\mathrm{obs}}}{\sigma_{\eta}}\right)^{2}.$ (4)
The latter term is a Gaussian constraint term for the background rate, with respect to the observed background rate (from phase I) and its uncertainty ($\eta_{\mathrm{obs}}$ and $\sigma_{\eta}$, respectively). $h_{4}(i)$ is the observed count rate model for phase IV in time bin ii. The expected counts in each time bin, $\nu(t_{i})$, is derived from the expected rate trend from phase III (see Eq. 1) multiplied by $R(t_{i})^{-1}$, plus $\eta$. Given the expected 15.1 cm average track length of 214Po events Ziegler:2010 relative to the 30 cm diameter of the SPC, some events may not deposit all their energy within the detector volume. Further, some fraction of 214Po isotopes are expected to settle on the inner surface of the detector before decaying, increasing the fraction of events that will not deposit their full energy in the SPC gas. This could result in an underestimation of the actual rate of 214Po. Consequently, the expected and observed 214Po event rates in phase IV are both conservatively reduced by 50% in the calculation of $R(t_{i})^{-1}$, increasing the statistical uncertainty in this result.
Figure 6: Top panel Expected decay rates with either the quality cut (light green curves) or with the addition of the 214Po cut applied (dark green curves), derived from modeling of the phase III data from each campaign (green histograms), as well as the phase IV data (red histograms). Bottom panel The 90% confidence level (CL) intervals of the inverse R-values over time, obtained with the quality cut (blue bands) and 214Po cut (50% reduced signal efficiency) additionally applied (red bands). The vertical dashed line defines the reference time where the extrapolated rate with quality cut (light green line) is equal to 76.5 Hz.
A $90\%$ confidence level (CL) interval is calculated for each time bin using the Feldman-Cousins method Feldman:1997qc . This process is repeated for 500 randomly–drawn MCMC samples from the fit of the phase III rate model and the phase IV observed rate model to include statistical uncertainty for both fits. This yields an ensemble of slightly–varying 90% CL intervals for $R(t_{i})^{-1}$ for each time bin, of which we take the median result for each. The $R(t_{i})^{-1}$ 90% CL intervals are shown in the bottom panels of Fig. 6 for each campaign. From these results, we determine the final R-value, $R(t)$, at the reference time, which is denoted by vertical dashed lines in Fig. 6. The derived values and their associated 90% CL intervals (upper and lower) for both cuts are summarized in Table 2. To quantitatively compare the R-value across all campaigns, the 90% lower confidence limit (LCL) of $R$ at the reference time is reported as the final result for each campaign and shown in Fig. 7 with or without the 214Po cut (with 50% reduced signal efficiency) applied.
@@ 4.2 Radon reduction ratio calculation
Table 2: 90% CL intervals (lower and upper limits) of R-values at the reference time with only the quality cut or also with the 214Po cut (reduced by 50%) applied for each campaign. The 90% lower confidence limit (LCL) is reported as the final result for each campaign and shown in Fig. 7
Campaign 1 Campaign 2 Campaign 3
Quality cut [4006.9, inf] [3803.3, inf] [5.4086, 5.5079]
@@
Figure 7: R-values at 90% lower confidence limit (LCL), obtained at the reference time, with the trap filled with silver-zeolite (campaigns 1 and 2) and activated charcoal (campaign 3). The blue and red bars represent the R-values obtained with either just the quality cut or also with the 214Po cut applied (with 50% reduced signal efficiency), respectively.
4.3 K-factor estimation
To enable a better comparison with previous studies 10.1093/ptep/ptad160 ; Heinitz:2023 , we also estimate the adsorption constant, or K-factor [m3/kg] obtained with our gas mixture 97% argon and 3% methane. With our current setup and available data, we cannot apply the standard definitions used in Ref. VESELSKA2026134640 . Instead, we follow the approach of Ref. Sone:2024gox , assuming that the radon concentration in our system—dominated by the detector volume—is proportional to the circulation time. In this framework, the R‑value is defined as the inverse of the radon concentration ratio (Eqs. 2–3 in Sone:2024gox ), which allows us to determine the retention time required to compute the estimation of the K-factor for our gas mixture.
With a detector volume of 14 L, a flow rate of 1 L/min, 10 g of adsorbent, and the 90% LCL R‑values with quality cuts from Table 2, we obtain K = 6.1720 m3/kg for activated charcoal (campaign 3), and K = 5608.3 and 5323.2 m3/kg for silver-zeolite Ag-ETS-10 from campaigns 1 and 2, respectively.
The K-factor obtained for activated charcoal is consistent with standard values at room temperature reported in the literature, typically in the range of K = 3–10 m3/kg VESELSKA2026134640 . For the Ag-ETS-10, a result of K = 5323.2–5608.3 m3/kg is comparable to those at room temperature reported in Ref. 10.1093/ptep/ptad160 (K = 1400 m3/kg) and Ref. Heinitz:2023 (K = 3400 m3/kg) in nitrogen gas and (K = 4300 m3/kg) in air Heinitz:2023 . However, it remains significantly lower than the value measured in pure argon (K $\approx$ 20 000 m3/kg) Heinitz:2023 .
It is important to note that differences in detector technology and operating pressure prevent a direct comparison. In addition, the K-factor strongly depends on the type of carrier gas, and no compatible measurements exist for our specific gas mixture or for the low operating pressure used here. Furthermore, the absence of detectable radon re‑emission in campaigns 1 and 2 with silver-zeolite (as observed in Ref. Heinitz:2023 for pure argon) suggests that the actual retention time may be significantly longer than our estimates. This would imply that our derived K-factor is likely conservative. Because our result is an estimation rather than a fully validated measurement, we intend to adapt our closed-loop system for future campaigns—particularly those planned at lower temperatures—to measure the radon activity directly inside the trap and to extend the post-trap-opening measurement period. These improvements will provide a more consistent method for determining the K-factor to compare with existing results.
5 Conclusion and outlook
A comparison of the silver-zeolite Ag-ETS-10 and activated charcoal campaigns reveals several critical insights. First, the silver-zeolite trap demonstrates significantly higher efficiency in radon removal, with R-values three orders of magnitude greater than those observed with the activated charcoal trap. This conclusion is compatible with the relative K‑factor estimations derived from alternative method Sone:2024gox , which reflect a similar contrast in performance between the two adsorbents. Additionally, the rate ratio, which compares event rates after the trap is opened to those before radon diffusion, approaches 1 for the silver-zeolite campaigns, demonstrating near-complete radon removal from the system. In contrast, the activated charcoal campaign exhibits residual radon levels in the rate ratio that are two to three orders of magnitude higher. Finally, in all measurements, the silver-zeolite campaigns show consistent performance, as evidenced by similar results across both campaigns, confirming reliable and reproducible outcomes.
We demonstrate that silver-zeolite Ag-ETS-10 achieves superior radon removal efficiency at room temperature compared to activated charcoal, which is typically used at cryogenic temperatures. Our innovative closed-loop system offers a novel approach beyond previous methods and is the first to demonstrate validated performance under realistic operating conditions, such as those present in dark-matter-type detectors. These findings not only confirm and extend previous studies Obrien:2021 ; Heinitz:2023 ; 10.1093/ptep/ptad160 ; Ogawa:2022tku ; Sone:2024gox ; 10.1093/ptep/ptaf007 , but also highlight the strong potential of silver-zeolite for ambient-temperature radon reduction systems in rare-event search experiments. This advancement is particularly important for underground laboratories, where elevated radon levels compromise the ultra-low background environments required for dark matter and neutrino physics experiments Cui:2023fcd ; Huang:2024 . Integrating silver-zeolite filters into ventilation systems and clean rooms can substantially reduce radon concentrations and enhance the supporting capabilities of these facilities Kamaha:2022ooe ; Baudis:2022qjb . Moreover, eliminating the need for cryogenic cooling offers significant operational advantages by simplifying system design and reducing the overall complexity of radon reduction systems for future ton-scale detectors and beyond. The development of room-temperature radon filters could enable useful applications across multiple sectors, including residential basements, healthcare facilities, and other specialized industrial environments. Since our experimental campaigns with silver-zeolite Ag-ETS-10 have been conducted with an actual dark matter-type detector, they serve as a proof of principle for its application in current and future rare event searches.
Acknowledgment
This work was supported by the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute and through the University of Alberta startup fund UOFAB Startup Piro. M.-C. Piro. also acknowledges funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) Subatomic Physics Discovery Grant (project) No. SAPPJ-2024-00025, from which P. Gros. and Y. Deng. were further supported. The work of D. Durnford. and C. Garrah. was also supported by the NSERC Graduate Scholarships Doctoral and Master programs, respectively. We are grateful for the support of Compute Ontario ontario , the BC DRI Group, and the Digital Research Alliance of Canada alliance required to carry out the simulation needed for this manuscript. We thank Hervé Le Provost (IRFU-CEA/Saclay) for his software development work in our data acquisition system. We also thank the NEWS-G collaboration for their technical support provided for this work and the review of this manuscript. We also thank Daniel Kuznicki for assisting with the silver-zeolite Ag-ETS-10 sample obtained from Extraordinary Adsorbents Inc. of Edmonton, Alberta ExtraEdmonton1 .